# Discovering modular solutions that generalize compositionally

### v1

@@ 3.3 Empirical verification in the theoretical setting
Identification requires connected support. We construct task distributions with both connected and disconnected support. Figure 2B shows that the student parameter modules have a high alignment both for a continuous and discrete task distribution when the task support is connected. Disconnecting the task support leads to a noticeable degradation.
Identification is sensitive to over-parameterization. Given a task distribution with compositional and connected support, we investigate the effect of over-parameterization on identification. Specifically, we vary the ratio of the number of hidden units $\hat{h}$ and modules $\hat{M}$ of the student with respect to the teacher. While Theorem 1 holds only for $\hat{h}=h$ and $M=\hat{M}$, optimizing the loss well enough typically requires slight over-parameterization. Figure 2C shows that a slight over-parameterization is indeed beneficial for linear identification. Yet, beyond $\hat{h}=8h$ and $\hat{M}=2M$, the identification noticeably degrades for discrete training tasks and cannot be alleviated by adjusting the weight decay (c.f. Table A1 in Appendix B.1). For a continuous task distribution, over-parameterization appears unproblematic, indicating that the learning dynamics introduce a beneficial inductive bias.
@@ 4.1 Hyperteacher
Figure 3: A In the hyperteacher we pick between 1 to $K$ of the $M$ teacher modules, adding them in parameter space to create a task. B ANIL and MAML fail to generalize to OOD tasks regardless of the support of the training distribution, while hypernetworks achieve good OOD accuracy only when the task support is compositional. C+D When the task distribution is compositional and connected the teacher modules can be linearly decoded from the student task embeddings. E Hypernetworks have high OOD accuracy for $K>1$ but low OOD accuracy for $K=1$. F OOD performance of hypernetworks is sensitive to over-parameterization in both the hidden dimension and module dimension. Error bars in B-F denote the standard error of the mean over 3 seeds.
@@
Compositional generalization is sensitive to over-parameterization. The theory in the discrete setting requires the student to have the same hidden and module dimension as the teacher. We test how sensitive the ability for compositional generalization in terms of OOD performance is to over-parameterization on these two axes. In Figure 3F, we observe that while a certain over-parameterization is necessary for the optimization to succeed, performance starts to decrease for larger over-parameterization.
@@ 5 Related work & Discussion
Our results are complementary to Lachapelle et al. (2023) demonstrating that L1-regularization enables disentangled identification in multi-task learning of sparse, linear feature combinations with a monolithic architecture. More generally, there has been widespread interest in assessing compositional generalization (Lake & Baroni, 2018; Ruis et al., 2020; Hupkes et al., 2020) as a crucial shortcoming of deep learning (Loula et al., 2018; Dessì & Baroni, 2019; Keysers et al., 2020) across various subfields (Battaglia et al., 2018; Atzmon et al., 2020; Montero et al., 2021; Liu et al., 2023). While with the advent of pretrained large language models, compositional abilities have seemingly come into reach (Zhou et al., 2023; Orhan, 2022; Furrer et al., 2021; Csordás et al., 2021) and many specialized architectures (Russin et al., 2019; Li et al., 2019; Gordon et al., 2020; Andreas, 2020; Liu et al., 2020; Lake, 2019; Nye et al., 2020; Kaiser & Sutskever, 2016) are outperformed by pretrained large language models (Furrer et al., 2021), it remains an open question whether the ability for compositional generalization extends beyond the training distribution (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023).
@@
Different from prior work we put particular emphasis on the multi-task setting and compositionality in weight space - a common and powerful motif in deep learning (Kumar & Daume III, 2012; Ruvolo & Eaton, 2013; Perez et al., 2018; Jayakumar et al., 2020). We show that linearly combining parameters can capture nonlinear latent structure raising an important question for future research on the interpretability of the resulting modules. We derive our theory in the infinite data regime while in practice we resort to meta-learning from finite samples. Indeed the sample complexity appears to scale unfavorably as the number of teacher modules increases (c.f. Figure A3). Running meta-learning on this scale is computationally expensive and future work would ideally amortize the inference procedure when scaling to larger problem instances (Zhmoginov et al., 2022). Moreover, while our identification result in the multi-task teacher-student setting indicates that over-parameterization of the student can be detrimental for compositional generalization, the situation is less clear outside the teacher-student setting necessitating further investigation. Finally, we note that we characterized the solutions of the modular teacher-student setting at equilibrium in the infinite data limit. Even when the modular solution constitutes a global optimum, we cannot provide any guarantees that it can be reached with gradient-based learning. Jarvis et al. (2023) have made an important first step towards analyzing the learning dynamics of modular and monolithic architectures using deep linear networks in synthetic, linearly-separable datasets. Understanding and analyzing the learning dynamics of the full modular teacher-student setting, as has previously been done for more restricted architectures (Seung et al., 1992; Saad & Solla, 1995; Goldt et al., 2019), remains an important open question, the answer to which might help elucidate why even modular architectures often collapse to monolithic solutions.
@@ Acknowledgements
Simon Schug Original idea, conceptualization of the project, development of theory, experimental design, implementation, running experiments, writing of the main text, writing of the appendix, literature review, creation of figures.
Seijin Kobayashi Original idea, conceptualization of the project, development of theory, proving main theorems, experimental design, implementation, running experiments, helped writing of the main text, writing of the appendix, helped creation of figures.
@@
Razvan Pascanu Senior project supervision, conceptualization of the project, provided feedback on the theory, provided feedback on experimental design, provided feedback in regular project meetings, provided feedback on the manuscript.
João Sacramento Senior project supervision, original idea, conceptualization of the project, helped develop theory, experimental design, provided feedback in regular project meetings, provided feedback on the manuscript.
Angelika Steger Senior project supervision, conceptualization of the project, development of theory, helped proving main theorems, provided feedback in regular project meetings, provided detailed feedback on the manuscript.
@@
To ensure the reproducibility of our work, we are providing the code for all our experiments as part of the submission and complement the main text with experimental details in Appendix C. For the theoretical results, we are complementing our intuitive exposition in the beginning of Section 3 and empirical verification in Section 3.3 with detailed proofs in Section A and example cases in Section A.2.2
@@ A.1 Single task parameter identification for ReLU MLP
This section presents rigidity results that allows us, under the right conditions, to identify the learned parameters of a student feed forward network learning to imitate the output of a teacher. Here the student may have the same or even a wider architecture. To our knowledge this is the first such results for ReLU two-layer MLP.
@@
  • no output weight $a_{k}$ is 0
  • no input weight vector ${\bm{w}}_{k}$ is 0
@@
A simple sufficient condition for the irreducibility of a ReLU MLP is that the teacher has no two input weights ${\bm{w}}_{k}$ that are colinear, and no output weight $a_{k}$ that is 0. While this is not a necessary condition, we will restrict our identification results to such teacher networks for simplicity, cf. Section A.1.2 for the general case.
@@ A.1.1 Identification with ReLU MLP
[…]
@@
then there exists a partition $\mathcal{S}_{1},\mathcal{S}_{2}$ of $[1..h]$ and a permutation $\sigma$ of the neurons in the hidden layer such that we have
@@
The idea of the proof consists in noticing that 0 loss means that a ReLU MLP constructed by concatenating the teacher and student hidden neurons need to be functionally 0 everywhere. By grouping together the teacher and student weight vectors that are colinear with each other, we then show that the sum of each of hidden neurons belonging to the same group need to be 0 or a linear function in $\bm{x}$}. Finally, because no two teacher weight vectors are colinear, we show that necessarily a student weight needs to be colinear with each teacher weight. The permutation $\sigma$ consists in this mapping, and the partition ${\cal S}_{1},{\cal S}_{2}$ in respectively student weights that are positively, resp. negatively colinear with the teacher.
@@
To now deduce some consequences from this equation we introduce some notation. For every ${\bm{w}}\in{\cal W}$ we denote by ${\cal C}_{\bm{w}}$ the set of all vectors in $\cal W$} that are collinear with $\bm{w}$} (including $\bm{w}$} itself) and with ${\cal N}_{\bm{w}}:={\cal W}\setminus{\cal C}_{\bm{w}}$ all vectors in $\cal W$} that are not collinear with $\bm{w}$}. Furthermore we use ${\cal C}_{\bm{w}}={\cal C}_{\bm{w}}^{+}\cup{\cal C}_{\bm{w}}^{-}$ to denote those vectors that are positively resp. negatively collinear with $\bm{w}$}.
Consider some arbitrary ${\bm{w}}^{\prime}\in{\cal W}$. By definition of ${\cal C}_{{\bm{w}}^{\prime}}$ and ${\cal N}_{{\bm{w}}^{\prime}}$ we can find a vector ${\bm{x}}_{0}\in\mathbb{R}^{n}$ such that
@@
[…]
@@
[…]
@@
[…]
@@
$\sum_{i\in{\cal S}_{2}}\hat{a}_{\sigma(i)}\hat{\bm{w}}_{\sigma(i)}{\bm{x}}\cdot{\mathbbm{1}}_{\hat{\bm{w}}_{\sigma(i)}{\bm{x}}<0}=-\sum_{i\in{\cal S}_{2}}a_{i}{\bm{w}}_{i}{\bm{x}}\cdot{\mathbbm{1}}_{{\bm{w}}_{i}{\bm{x}}<0}.$
@@
Under the assumptions from Theorem 4, but with $\hat{h}\geq h$, there exists a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ of $1..\hat{h}$] for some $h^{\prime}$such that we have
@@
Firstly, we denote by ${\cal S}_{0}$ the indices of student vectors such that $\hat{a}\hat{\bm{w}}=0$. They correspond to student neurons that have 0 activities everywhere. This gives equation 21. Without loss of generality, we will exclude all student vectors in ${\cal S}_{0}$ from $\cal W$}, and redefine $\cal C$}, etc accordingly.
Secondly, recall that we assumed that no two rows in the teacher network are colinear and that in the teacher network no weight in $\bm{a}$} is zero. That is, for any ${\bm{w}}^{\prime}\in{\cal W}$ from the teacher network ${\cal C}_{{\bm{w}}^{\prime}}^{+}\cup{\cal C}_{{\bm{w}}^{\prime}}^{-}$ needs to contain at least one vector from the student network. For all teacher vector ${\bm{w}}_{i}$ where $i\in[1..h]$ we denote by resp. ${\cal S}_{1}(i),{\cal S}_{2}(i)$ the subsets of $1..\hat{h}$] corresponding to indices of student vectors that belong to ${\cal C}_{{\bm{w}}_{i}}^{+}$ resp. ${\cal C}_{{\bm{w}}_{i}}^{-}$. Observe that we can then rewrite equation 12 as
@@
Clearly, ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ form a partition of $1..\hat{h}$].

### v2

@@ 3.3 Empirical verification in the theoretical setting
Identification requires connected support. We construct task distributions with both connected and disconnected support. Figure 2C shows that the student parameter modules have a high alignment both for a continuous (${\bm{z}}\in{\mathbb{R}}^{m}$) and discrete (${\bm{z}}\in\{0,1\}^{m}$) task distribution when the task support is connected. Disconnecting the task support leads to a noticeable degradation.
Identification is sensitive to overparameterization. Given a task distribution with compositional and connected support, we investigate the effect of overparameterization on identification. Specifically, we vary the ratio of the number of hidden units $\hat{h}$ and modules $\hat{M}$ of the student with respect to the teacher. While Theorem 1 holds only for $\hat{h}=h$ and $M=\hat{M}$, optimizing the loss well enough typically requires slight overparameterization. Figure 2D shows that a slight overparameterization is indeed beneficial for linear identification. Yet, beyond $\hat{h}=8h$ and $\hat{M}=2M$, the identification noticeably degrades for discrete training tasks and cannot be alleviated by adjusting the weight decay (c.f. Table A1 in Appendix B.1). For a continuous task distribution, overparameterization appears unproblematic, indicating that the learning dynamics introduce a beneficial inductive bias.
@@ 4.1 Hyperteacher
Figure 3: A In the hyperteacher we pick between 1 to $K$ of the $M$ teacher modules, adding them in parameter space to create a task. B ANIL and MAML fail to generalize to OOD tasks regardless of the support of the training distribution, while hypernetworks achieve good OOD accuracy only when the task support is compositional. C+D When the task distribution is compositional and connected the teacher modules can be linearly decoded from the student task embeddings. E Hypernetworks have high OOD accuracy for $K>1$ but low OOD accuracy for $K=1$. F OOD performance of hypernetworks is sensitive to overparameterization in both the hidden dimension and module dimension. Error bars in B-F denote the standard error of the mean over 3 seeds.
@@
Compositional generalization is sensitive to overparameterization. The theory in the discrete setting requires the student to have the same hidden and module dimension as the teacher. We test how sensitive the ability for compositional generalization in terms of OOD performance is to overparameterization on these two axes. In Figure 3F, we observe that while a certain overparameterization is necessary for the optimization to succeed, performance starts to decrease for larger overparameterization.
@@ 5 Related work & Discussion
Our results complement Lachapelle et al. (2023) who demonstrate that L1-regularization enables disentangled identification in multi-task learning of sparse, linear feature combinations. Generally, there has been widespread interest in assessing compositional generalization (Lake & Baroni, 2018; Ruis et al., 2020; Hupkes et al., 2020) as a crucial shortcoming of deep learning (Loula et al., 2018; Dessì & Baroni, 2019; Keysers et al., 2020) across various subfields (Battaglia et al., 2018; Atzmon et al., 2020; Montero et al., 2021; Liu et al., 2023). While with the advent of pretrained large language models, compositional abilities have seemingly come into reach (Zhou et al., 2023; Orhan, 2022; Furrer et al., 2021; Csordás et al., 2021) and many specialized architectures (Russin et al., 2019; Li et al., 2019; Gordon et al., 2020; Andreas, 2020; Liu et al., 2020; Lake, 2019; Nye et al., 2020; Kaiser & Sutskever, 2016) are outperformed by pretrained large language models (Furrer et al., 2021), it remains an open question whether the ability for compositional generalization extends beyond the training distribution (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023).
@@
Different from prior work we put particular emphasis on the multi-task setting and compositionality in weight space - a common and powerful motif in deep learning (Kumar & Daume III, 2012; Ruvolo & Eaton, 2013; Perez et al., 2018; Jayakumar et al., 2020; Dimitriadis et al., 2023). We show that linearly combining parameters can capture nonlinear latent structure raising an important question for future research on the interpretability of the resulting modules. We derive our theory in the infinite data regime while in practice we resort to meta-learning from finite samples. Indeed the sample complexity appears to scale unfavorably as the number of teacher modules increases (c.f. Figure A3). Running meta-learning on this scale is computationally expensive and future work would ideally amortize the inference procedure when scaling to larger problem instances (Zhmoginov et al., 2022). Moreover, while our identification result in the multi-task teacher-student setting indicates that overparameterization of the student can be detrimental for compositional generalization, the situation is less clear outside the teacher-student setting necessitating further investigation. Finally, we note that we characterized the solutions of the modular teacher-student setting at equilibrium in the infinite data limit. Even when the modular solution constitutes a global optimum, we cannot provide any guarantees that it can be reached with gradient-based learning. Jarvis et al. (2023) have made an important first step towards analyzing the learning dynamics of modular and monolithic architectures using deep linear networks in synthetic, linearly-separable datasets. Understanding and analyzing the learning dynamics of the full modular teacher-student setting, as has previously been done for more restricted architectures (Seung et al., 1992; Saad & Solla, 1995; Goldt et al., 2019), remains an important open question, the answer to which might help elucidate why even modular architectures often collapse to monolithic solutions.
@@ Acknowledgements
Simon $*{}^{*}$ Original idea, conceptualization of the project, development of theory, experimental design, implementation, running experiments, writing of the main text, writing of the appendix, literature review, creation of figures.
Seijin $*{}^{*}$ Original idea, conceptualization of the project, development of theory, proving main theorems, experimental design, implementation, running experiments, helped writing of the main text, writing of the appendix, helped creation of figures.
@@
Razvan ${}^{{\dagger}}$ Senior project supervision, conceptualization of the project, provided feedback on the theory, provided feedback on experimental design, provided feedback in regular project meetings, provided feedback on the manuscript.
João ${}^{{\dagger}}$ Senior project supervision, original idea, conceptualization of the project, helped develop theory, experimental design, provided feedback in regular project meetings, provided feedback on the manuscript.
Angelika ${}^{{\dagger}}$ Senior project supervision, conceptualization of the project, development of theory, helped proving main theorems, provided feedback in regular project meetings, provided detailed feedback on the manuscript.
@@
To ensure the reproducibility of our work, we are providing the code for all our experiments online at https://github.com/smonsays/modular-hyperteacher and complement the main text with experimental details in Appendix C. For the theoretical results, we are complementing our intuitive exposition in the beginning of Section 3 and empirical verification in Section 3.3 with detailed proofs in Section A and example failure cases to build intuition in Section A.3.
@@ A.1 Single task parameter identification for ReLU MLP
This section presents rigidity results that allows us, under the right conditions, to identify the learned parameters of a student MLP learning to imitate the output of a teacher. Here the student may have the same or even a wider architecture. To our knowledge this is the first such results for ReLU two-layer MLP.
@@
  • no output weight $a_{k}$ is 00
  • no input weight vector ${\bm{w}}_{k}$ is 00
@@
A simple sufficient condition for the irreducibility of a ReLU MLP is that the teacher has no two input weights ${\bm{w}}_{k}$ that are colinear, and no output weight $a_{k}$ that is 00. While this is not a necessary condition, we will restrict our identification results to such teacher networks for simplicity, cf. Section A.1.2 for the general case.
@@ A.1.1 Identification with ReLU MLP
Assume a teacher and student two-layer network parameterized by resp. $({\bm{W}},{\bm{a}})$ and $(\hat{{\bm{W}}},\hat{{\bm{a}}})$, of hidden dimensions $-^h=\hat{h}$ and a single output unit such that
@@
[…]
@@
The idea of the proof consists in noticing that 00 loss means that a ReLU MLP constructed by concatenating the teacher and student hidden neurons need to be functionally 00 everywhere. By grouping together the teacher and student weight vectors that are colinear with each other, we then show that the sum of each of hidden neurons belonging to the same group need to be 00 or a linear function in ${\bm{x}}$. Finally, because no two teacher weight vectors are colinear, we show that necessarily a student weight needs to be colinear with each teacher weight. The permutation $\sigma$ consists in this mapping, and the partition ${\cal S}_{1},{\cal S}_{2}$ in respectively student weights that are positively, resp. negatively colinear with the teacher.
@@
To now deduce some consequences from this equation we introduce some notation. For every ${\bm{w}}\in{\cal W}$ we denote by ${\cal C}_{\bm{w}}$ the set of all vectors in ${\cal W}$ that are collinear with ${\bm{w}}$ (including ${\bm{w}}$ itself) and with ${\cal N}_{\bm{w}}:={\cal W}\setminus{\cal C}_{\bm{w}}$ all vectors in ${\cal W}$ that are not collinear with ${\bm{w}}$. Furthermore we use ${\cal C}_{\bm{w}}={\cal C}_{\bm{w}}^{+}\cup{\cal C}_{\bm{w}}^{-}$ to denote those vectors that are positively resp. negatively collinear with ${\bm{w}}$. Consider some arbitrary ${\bm{w}}^{\prime}\in{\cal W}$. By definition of ${\cal C}_{{\bm{w}}^{\prime}}$ and ${\cal N}_{{\bm{w}}^{\prime}}$ we can find a vector ${\bm{x}}_{0}\in\mathbb{R}^{n}$ such that
@@
and an open ball $B$ around ${\bm{x}}_{0}$ such that we have for all ${\bm{w}}\in{\cal N}_{{\bm{w}}^{\prime}}$: ${\bm{w}}^{\top}{\bm{x}}\neq 0$ for all ${\bm{x}}\in B$. Using ${\mathbbm{1}}_{\cal E}$ to denote the indicator for the event ${\cal E}$, i.e. ${\mathbbm{1}}_{\cal E}\equiv 1$ if ${\cal E}$ holds and ${\mathbbm{1}}_{\cal E}\equiv 0$ otherwise, we thus have
@@
$0=(\sum_{{\bm{w}}\in{\cal N}_{{\bm{w}}^{\prime}}}c_{\bm{w}}{\mathbbm{1}}_{{\bm{w}}^{\top}{\bm{x}}_{0}>0}\cdot{\bm{w}}+\sum_{{\bm{w}}\in{\cal C}_{{\bm{w}}^{\prime}}^{+}}c_{\bm{w}}{\bm{w}})^{\top}{\bm{x}}\qquad\text{ for all ${\bm{x}}\in B^{+}$}$
@@
$0=(\sum_{{\bm{w}}\in{\cal N}_{{\bm{w}}^{\prime}}}c_{\bm{w}}{\mathbbm{1}}_{{\bm{w}}^{\top}{\bm{x}}_{0}>0}\cdot{\bm{w}}+\sum_{w\in{\cal C}_{{\bm{w}}^{\prime}}^{-}}c_{\bm{w}}{\bm{w}})^{\top}{\bm{x}}\qquad\text{ for all ${\bm{x}}\in B^{-}$}$
@@
$-\sum_{i\in{\cal S}_{2}}\hat{a}_{\sigma(i)}\hat{\bm{w}}_{\sigma(i)}{\bm{x}}\cdot{\mathbbm{1}}_{\hat{\bm{w}}_{\sigma(i)}{\bm{x}}<0}=-\sum_{i\in{\cal S}_{2}}a_{i}{\bm{w}}_{i}{\bm{x}}\cdot{\mathbbm{1}}_{{\bm{w}}_{i}{\bm{x}}<0}.$
@@
Under the assumptions from Theorem 4, but with $-^\hat{h}\geq h$, there exists a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ of $[1..\hat{h}]$ for some $h^{\prime}$such that we have
@@
Firstly, we denote by ${\cal S}_{0}$ the indices of student vectors such that $\hat{a}\hat{\bm{w}}=0$. They correspond to student neurons that have 00 activities everywhere. This gives equation 21. Without loss of generality, we will exclude all student vectors in ${\cal S}_{0}$ from ${\cal W}$, and redefine ${\cal C}$, etc accordingly.
Secondly, recall that we assumed that no two rows in the teacher network are colinear and that in the teacher network no weight in ${\bm{a}}$ is zero. That is, for any ${\bm{w}}^{\prime}\in{\cal W}$ from the teacher network ${\cal C}_{{\bm{w}}^{\prime}}^{+}\cup{\cal C}_{{\bm{w}}^{\prime}}^{-}$ needs to contain at least one vector from the student network. For all teacher vector ${\bm{w}}_{i}$ where $i\in[1..h]$ we denote by resp. ${\cal S}_{1}(i),{\cal S}_{2}(i)$ the subsets of $[1..\hat{h}]$ corresponding to indices of student vectors that belong to ${\cal C}_{{\bm{w}}_{i}}^{+}$ resp. ${\cal C}_{{\bm{w}}_{i}}^{-}$. Observe that we can then rewrite equation 12 as
@@
Clearly, ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ form a partition of $[1..\hat{h}]$.